OpenStudy (anonymous):

Equation I: y = 3x + 5 Equation II: x + y = 3 Solve the above system of linear equations by using the substitution method.

OpenStudy (anonymous):

x + 3x + 5 = 3 4x + 5 = 3 4x = -2 x = -2/4 = -1/2 y = 7/2